OpenStudy (anonymous):

log3 (5) +log3 ( x2+4) =3

OpenStudy (solomonzelman):

\[\log_3 (5) +\log_3 ( x^2+4) =3\]\[\log_3 (5) +\log_3 ( x^2+4) =3 \times \log_33\]\[\log_3 (5) +\log_3 ( x^2+4) =\log_33^3\]\[\log_3 (5) +\log_3 ( x^2+4) =\log_327\]Then apply,\[logA+logB=\log(AB)\]and solve for x

OpenStudy (anonymous):

Thank You! But now how do I solve for x?

OpenStudy (solomonzelman):

tell me what do you get after applying \[\log(a)+\log(b)=\log(a \times b)\]

OpenStudy (anonymous):

I dont know what a and b are

OpenStudy (solomonzelman):

in your case, a is 5 and b is x²+4

OpenStudy (anonymous):

So then I should get log (5) +log (x²+4) = log (5)(x²+4)

OpenStudy (solomonzelman):

Yes, \[\log_3[~~5(x^2+4)~~~]\]Can you expand the parenthesis inside the log ?

OpenStudy (anonymous):

I think you get log3 [5x2 + 20]

OpenStudy (solomonzelman):

yes, correct.

OpenStudy (solomonzelman):

I hate disconnecting !

OpenStudy (solomonzelman):

Made a mistake, sorry. this time it will be correct. \(\LARGE\color{blue}{ \bf log_3\color{red} { 5x^2+20 } = log_3\color{red} { 27} }\) Do you see what you can remove/imply ?